OpenStudy (anonymous):

Thomas surveyed his class to find the number of pets that each student had. He complied his results into a probability distribution which is shown. What is the expected number of pets for a student in his class?

OpenStudy (anonymous):

Here is the probability distribution that goes along with it http://screencast.com/t/LGJmj0Wl7Pn

OpenStudy (anonymous):

So far i have 0(.04) + .3(1) + .25(2) + .005(3) = .95, However i can't remember what operation to use to find the expected outcome of all students.

OpenStudy (amistre64):

that 3 or more part has me overthinking it ... but otherwise your process is fine, \[E(X)=\sum x~ p(x)\]

OpenStudy (anonymous):

well the asnwer doesnt show .95 its just ~1, ~2, ~3, ~4

OpenStudy (anonymous):

unless they want me to round up?

OpenStudy (amistre64):

.95 is approx 1 isnt it?

OpenStudy (anonymous):

yes here ill see if it is correct

OpenStudy (anonymous):

yes it was

OpenStudy (anonymous):

Would you mind helping me with 1 more question?

OpenStudy (amistre64):

i spose they wanted a more realistic value ... kind of hard to have .95 pets :)

OpenStudy (amistre64):

i can try

OpenStudy (anonymous):

here it is http://screencast.com/t/w08FenBTPr

OpenStudy (anonymous):

is the average 80%?

OpenStudy (anonymous):

100+80+60+0 = 240 240/4 =60, well 60%

OpenStudy (amistre64):

what is the probability of each outcome?

OpenStudy (amistre64):

.5^2 ...

OpenStudy (amistre64):

P(0 correct) = (3 0) .5^0 .5^3 P(1 correct) = (3 1) .5^1 .5^2 P(2 correct) = (3 2) .5^2 .5^1 P(3 correct) = (3 3) .5^3 .5^0

OpenStudy (amistre64):

E(X)= 0 P(0)+ 1P(1)+ 2P(2)+ 3P(3) you agree?

OpenStudy (amistre64):

hmm, maybe better as 0 60 80 and 100 .... not 0 1 2 3

OpenStudy (amistre64):

or, find expected number of points and determine it from the grading rubric

OpenStudy (anonymous):

how are you getting 5^2, im getting 5.2

OpenStudy (anonymous):

i just took 3(.1)+2(.8)+1(.6)

OpenStudy (amistre64):

each question is a 50/50 chance of being right, so for any given test (this is a binomial approach) we have: rrr rrw rwr wrr rww wrw wwr www

OpenStudy (amistre64):

0 right = 3 wrong there is 1 www, and w = .5 1 right is 3 cases, rww, wrw, wwr, r=.5 and w=.5 etc

OpenStudy (amistre64):

.5^3 is the result that the probability of right = probability of wrong when guessing between 2 options. .5 and .5

OpenStudy (amistre64):

does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then each permutation result has a .5^3 probability 0 correct = 1 case of .5^3 1 correct = 3 cases of .5^3 2 correct = 3 cases of .5^3 3 correct = 1 case of .5^3 it matter not if we use the number of points scored or the number of questions correct as a basis .... they have equal weight regardless, but you might feel safer finding the expected (mean) of number correct.

OpenStudy (amistre64):

0(1*.5^3)+1(3*.5^3)+2(3*.5^3)+3(1*.5^3) 3*.5^3+6*.5^3+3*.5^3 .5^3(3+6+3) 12*.5^3

OpenStudy (anonymous):

i was just doing that... in my calculator lol

OpenStudy (anonymous):

i just get 1.5

OpenStudy (amistre64):

so our average number correct is 1.5 ... rnd to 2?

OpenStudy (anonymous):

yep

OpenStudy (amistre64):

try it using point values instead and you should get the same result but in point format instead of number correct format

OpenStudy (amistre64):

0(1*.5^3)+60(3*.5^3)+80(3*.5^3)+100(1*.5^3) we get 65 as the average points earned for guessing

OpenStudy (amistre64):

so do we round to 80 or stay at 60 ... but thats up to how your material wants you to proceed

OpenStudy (anonymous):

just stay but here

OpenStudy (anonymous):

There are eight possible ways that the questions on the quiz can be answered. If R stands for a right answer and W stands for a wrong answer, the eight ways are: RRR, WRR, RWR, RRW, WWR, WRW, RWW, WWW. This means that the probability of 3 right answers is 18 , the probability of 2 right answers is 38 , the probability of 1 right answer is 38 , and the probability of 0 right answers is 18 . To find the average number of points a student is likely to be awarded on the quiz, find the expected value:(18)(100) + (38)(80) + (38)(60) + (18)(0) = 65.

OpenStudy (anonymous):

thanks for the help, relly appreciate it. :P gave you best

OpenStudy (amistre64):

youre welcome :)